Papst Company is preparing its cash budget for the month of May. The following information is available concerning its accounts receivable (based on sales made to customers on open account): Actual credit sales for March $ 225,000 Actual credit sales for April $ 293,000 Estimated credit sales for May $ 438,000 Estimated collections in the month of sale 25 % Estimated collections in the first month after the month of sale 60 % Estimated collections in the second month after the month of sale 10 % Estimated provision for bad debts (made in the month of sale) 5 % The firm writes off all uncollectible accounts at the end of the second month after the month of sale.

Required: Determine for Papst Company for the month of May: 1. The estimated cash receipts from accounts receivable collections. 2. The gross amount of accounts receivable at the end of the month (after appropriate write-off of uncollectible accounts). 3. The net amount of accounts receivable at the end of the month. 4. Recalculate requirements 1 and 2 under the assumption that estimated collections in the month of sale equal 60% and in the first month following the month of sale equal 25%.

1. Estimated cash receipts?

2. Gross accounts receivable

3. Net accounts receivable

4a. Estimated cash receipts

4b. Gross accounts receivable
Answer #1

SOLUTION

1. Cash receipts from Accounts Receivable collections = 25% of May Sales + 60% of April Sales + 10% of March Sales

= (25%*$438,000) + (60%*293,000) + (10%*225,000)

= 109,500 + 175,800 + 22,500

= 307,800

2. Gross Account Receivable at the end of May = 75% of May Sales + 15% of April Sales

= ($438,000*75%) + ($293,000*15%)

= 328,500 + 43,950 = 372,450

3. Allowance for Doubtful Accounts = 5%*April Sales + 5% of May Sales

= ($293,000*5%) + ($438,000*5%)

= 14,650 + 21,900 = 36,550

Net Accounts Receivable at the end of May = Gross Account Receivable at the end of May - Allowance for Doubtful Accounts

= 372,450 - 36,550 = 335,900

4A. Cash receipts from Accounts Receivable collections = 60% of May Sales + 25% of April Sales + 10% of March Sales

= (60%*$438,000) + (25%*293,000) + (10%*225,000)

= 262,800 + 73,250 + 22,500

= 358,550

4B. Gross Account Receivable at the end of May = 40% of May Sales + 15% of April Sales

= (40%*$438,000) + (15%*293,000)

= 175,200 + 43,950

= 219,150